### v2

@@ 2.3.2 Vector magnetic gradiometer
Figure 3: Magnetic gradiometer sensitivity. The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and two transverse magnetic fields. The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer. The inset displays the frequency response of $B_{x}$ and $B_{y}$.
3 Discussion
@@
In summary, we introduce a FRF vector geomagnetic magnetometer. This system is achieved by applying a rotating field to a scalar atomic magnetometer, enabling it to fully determine the magnetic field vector. While the modulation slightly degrades the scalar performance from 40 $\mathrm{fT/\sqrt{Hz}}$ to 50 $\mathrm{fT/\sqrt{Hz}}$, it provides two additional polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$. This enhancement allows the vector magnetometer to precisely measure two transverse polar angles, eliminating the static heading error common in scalar atomic magnetometers. The unique vector axes are defined by the rotation plane of the applied field, increasing stability compared to relying on mechanical coil orthogonality.
@@
We provide a comprehensive study of the systematics and develop cosine-altering rotating field modulation to cancel out these effects. Benefits from frequency measurement, such vector magnetometers can achieve high fractional resolution. Additionally, the vector magnetometer retains the accuracy and metrological advantages of scalar atomic magnetometers, such as inherent calibration.
@@
4 Methods
4.1 Noise limit
Based on the CRLB [42] and the photon shot noise of the balanced polarimeter detection [43], the power spectral density $\rho(\omega)$ has a minimum when the measurement time $t\approx 2T_{2}$ (check Sec. III of supplemental material for more details). The total magnetic field measurement sensitivity is given by
$\delta B_{tot}\geq\frac{4}{\gamma kT_{2}\sqrt{\Phi_{pr}}}.$ (10)
where $k=lr_{e}cfnD(\nu)/2$, $l$ is the length of probe beam travel through the cell, $r_{e}$ is the classical electron radius, $f$ is the typical oscillator strength of the D line transition, $D(\nu)=(\nu-\nu_{0})/[(\nu-\nu_{0})^{2}+(\Delta\nu/2)^{2}]$, where $\Delta\nu$ is the optical full width at half maximum (FWHM), $\nu_{0}$ is the frequency of the D line transition. $T_{2}$ is the transverse relaxation time. $\Phi_{pr}$ is the total flux of photons of the probe beam per unit of time. If we keep the measurement time the same as the optimized measurement time of the total magnetic field $t=2T_{2}$, then the transverse magnetic field measurement sensitivity is given by
@@ 4.1 Noise limit
The derived fundamental sensitivities for measuring the scalar magnetic field and transverse magnetic field are based on the assumption that the fitting of these fields is independent. However, this assumption is not accurate when the modulation frequency is not high enough. An analytical solution for the sensitivities can be obtained by fitting the scalar and transverse magnetic fields simultaneously. This can be achieved using the Minimum-Variance Unbiased (MVU) estimator, similar to the approach described in Sections III and IV of the supplemental material. The results are more complex, but it is found that when the modulation frequency is fast enough, the assumption of independent fitting holds. ($\omega_{m}\gtrsim\pi/T_{2}$), the fundamental sensitives are almost the same as considering them independently.
For comparison, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers can be described as (Please refer to Sec. IV of the supplemental material for details.)
$\delta B_{tran\_SM}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (12)
If we assume $\omega_{m}=\pi/T_{2}$, the sensitivity of the transverse magnetic field of the FRF vector magnetometer can be written as
$\delta B_{tran\_FRF}=\frac{2\sqrt{2}\pi}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}$ (13)
The FRF vector magnetometer demonstrates a sensitivity in measuring transverse magnetic fields that is comparable to sequential modulation vector magnetometers. Additionally, it offers a higher bandwidth and is less susceptible to systematic effects.
@@ Acknowledgements
Author contributions statement
T.W. prepared the setup and conducted the experiments for the FRF vector magnetometers, proposed and implemented the CARFM modulation, modeled the spin evolution using rotation matrices and density matrices, derived the fundamental limits, analyzed the results, and prepared the manuscript. M.R. supervised the setup and experiments, modeled the spin evolution using rotation matrices, and discussed the CARFM modulation and dynamic heading error. W.L. assisted with modeling the static heading error and analyzing the polarization of the atoms. M.L., T.K., and E.F. provided the scalar magnetometer sensors and offered some valuable experimental advice. All authors reviewed the manuscript.
@@
Additional information
The corresponding author is responsible for submitting a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.

### v3

@@ II.3.2 Results of magnetic field sensitivities
Figure 3: Magnetic gradiometer sensitivity: The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields. The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer. The inset shows the frequency response for $B_{x}$ and $B_{y}$.
III Discussion
@@
In summary, we present a FRF vector geomagnetic magnetometer. This system is achieved by applying a rotating field to a scalar atomic magnetometer, enabling it to fully determine the magnetic field vector. While the modulation slightly degrades the scalar performance from 28 $\mathrm{fT/\sqrt{Hz}}$ to 35 $\mathrm{fT/\sqrt{Hz}}$, it provides two additional polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$. This enhancement allows the vector magnetometer to precisely measure two transverse polar angles, eliminating the static heading error common in scalar atomic magnetometers. The unique vector axes are defined by the rotation plane of the applied field, increasing stability compared to relying on mechanical coil orthogonality.
@@
We provide a comprehensive study of the systematics and develop peak-altering fast rotating field modulation to cancel out these effects. Benefits from frequency measurement, such vector magnetometers can achieve high fractional resolution. Additionally, the vector magnetometer retains the accuracy and metrological advantages of scalar atomic magnetometers, such as inherent calibration.
@@
IV Methods
IV.1 Shot noise limit
To derive the shot noise limit for the FRF magnetometer, we utilize the CRLB [41] along with photon shot noise considerations specific to balanced polarimeter detection [42]. Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Sec. IV of the supplemental material). For total magnetic field measurements, the sensitivity is expressed as
$\delta B_{tot}\geq\frac{4}{\gamma kT_{2}\sqrt{\Phi_{pr}}},$ (11)
where $k=lr_{e}cfnD(\nu)/2$, with $l$ being the probe beam path length through the cell, $r_{e}$ the classical electron radius, $f$ the typical oscillator strength of the D-line transition, and $D(\nu)=(\nu-\nu_{0})/[(\nu-\nu_{0})^{2}+(\Delta\nu/2)^{2}]$, where $\Delta\nu$ is the optical full width at half maximum (FWHM) and $\nu_{0}$ is the D-line transition frequency. Here, $T_{2}$ is the transverse relaxation time, and $\Phi_{pr}$ denotes the photon flux per unit time. This noise limit also applies to scalar atomic magnetometers when analyzing the free induction decay (FID) signal.
If the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
@@ IV.1 Shot noise limit
These derived sensitivities assume independent fitting of scalar and transverse magnetic fields. However, this assumption becomes inaccurate if the modulation frequency is not sufficiently high. Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Sections III and IV of the supplemental material. The results show that when modulation frequency is sufficiently fast ($\omega_{m}\gtrsim\pi/T_{2}$), independent fitting yields comparable sensitivities.
Based on experimental parameters, we estimate the shot noise sensitivities of the FRF magnetometer to be $\delta B_{tot}=0.04~{}\mathrm{fT/\sqrt{Hz}}$ and $\delta B_{tran}=0.8~{}\mathrm{fT/\sqrt{Hz}}$. However, experimental measurements show sensitivities worse than these theoretical limits, primarily due to magnetic noise from current sources.
For instance, an SRS DC205 supplies approximately 650 mA to generate the primary magnetic field along the z-axis, simulating a 50 $\mu$T Earth field. Despite modifications to reduce current noise, it remains a primary noise source, limiting overall sensitivity.
To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Sec. VI of the supplemental material for details):
$\delta B_{tran\_SM}=\frac{4\sqrt{2}}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}.$ (13)
If $\omega_{m}=\pi/T_{2}$, the sensitivity for transverse magnetic fields in the FRF vector magnetometer can be expressed as
$\delta B_{tran\_FRF}=\frac{2\sqrt{2}\pi}{\gamma kT_{2}\sin\theta\sqrt{\Phi_{pr}}}.$ (14)
Thus, the FRF vector magnetometer achieves transverse field sensitivity comparable to that of sequential modulation vector magnetometers while offering a higher bandwidth and reduced susceptibility to systematic effects.
@@ Acknowledgements
Author Contributions
T.W. prepared the setup and conducted experiments for the FRF vector magnetometers, proposed and implemented the modulation, modeled spin evolution using rotation and density matrices, derived fundamental limits, analyzed results, and prepared the manuscript. M.R. supervised the setup and experiments, contributed to spin evolution modeling with rotation matrices, and discussed modulation and dynamic heading error. W.L. assisted with modeling static heading error and analyzing atomic polarization. M.L., T.K., and E.F. provided scalar magnetometer sensors, offered valuable experimental advice, and provided constructive comments for manuscript revisions. All authors reviewed the manuscript.